# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ 6. Quantum polymorphisms of odd cycles
If $x=z$ the result is trivial so we may assume that $x\neq z$. We know from Corollary 30 that there exists $d\in\mathbb{N}$ and $h_{1},\dots h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$. Using eq. 1, we can write
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{\top}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{\top}\right)$
$=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}.$
@@
$Q_{x,a}Q_{y,b}Q_{z,c}=Q_{x,a}Q_{y,b}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{\top}=Q_{y,b}Q_{z,c},$
@@
Next, we focus on quantum polymorphisms of arity at least two. Henceforth in this section, we fix some arity $n\geq 2$. Furthermore, whenever a plus or minus sign is applied to an element of $\mathbb{Z}_{m}$ or to a tuple of such elements, it shall be meant modulo $m$.
@@
Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are Hermitian, it follows that
@@
We now apply the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
@@
In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is a Hermitian matrix, so
@@
We can now finalise the proof of the main result of this section.
@@
(10) $Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
@@
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}$
@@
where the second equality is eq. 10. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required.
Corollary 34.
Let $m\geq 3$ be an odd natural number. Then $\operatorname{qPol}(\mathbb{C}_{m})=\operatorname{\oplus Pol}(\mathbb{C}_{m})$. In particular, $\mathbb{C}_{m}$ has a commutativity gadget and $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable.
@@
$\mathbb{C}_{m}$ has a commutativity gadget by Theorem 29 and Theorem 16.
Note that $\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a simple path of length m2m-2. Since $\mathbb{C}_{m}$ has a commutativity gadget, it follows that $\mathbb{K}_{m}$ is q-definable in $\mathbb{C}_{m}$ by Proposition 13. By Proposition 10, $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ reduces to $\operatorname{CSP^{*}}(\mathbb{C}_{m})$, and $\operatorname{CSP^{*}}(\mathbb{K}_{m})$ is undecidable by Corollary 24. Thus, $\operatorname{CSP^{*}}(\mathbb{C}_{m})$ is undecidable. ∎
We conclude with the proof that quantum homomorphisms between different odd cycles can be contextual.
Proposition 35.
There is a quantum homomorphism $\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
Proof.
Consider for this proof $\mathbb{C}_{5}$ to have domain $\{0,1,2,5,6\}$ with edges between adjacent numbers, and $\mathbb{C}_{7}$ to have domain $[7]$. Consider the two classical endomorphisms $h,g\colon\mathbb{C}_{5}\to\mathbb{C}_{5}$ such that $h(2)=6,h(5)=0$ and $g(5)=0,g(2)=1$. Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{q}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space. Extend $Q$ by setting $Q_{3,a}=Q_{5,a}$ for $a\in\{0,1,2,5,6\}$. Choose a basis in which all the projectors are diagonal, and in this basis define $Q_{4,1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ and $Q_{4,6}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$. Then $Q\colon\mathbb{C}_{7}\to_{q}\mathbb{C}_{5}$ and $Q_{4,0},Q_{2,1}]\neq 0$. ∎
7. The non-oracular setting
In this section, we consider the non-oracular version of the results in Section 3. A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{qno}}\mathbb{B}$, is a quantum function $Q\colon A\to_{q}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)).
The following is the non-oracular version of the entangled CSP.
Definition 36 (Non-oracular entangled CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP^{**}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{qno}}\mathbb{A}$.
We now define quantum polymorphisms in the non-oracular setting.
@@ 7. The non-oracular setting
Let $\mathbb{A}$ be a relational structure. A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We denote by $\operatorname{qPol^{no}}(\mathbb{A})$ the set of non-oracular quantum polymorphisms of $\mathbb{A}$.
We will obtain a quantisation of the Galois connection between classical polymorphisms and relations. The following is the main result of this section.
@@
Then $\operatorname{CSP^{*}}(\mathbb{B})$ reduces to $\operatorname{CSP^{**}}(\mathbb{A})$ in logspace.
@@
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $\mathbb{A}_{S}$, and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A non-oracular q-definition for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
@@ 1. $(\mathbf{noq_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{q}\mathbb{A}_{S}$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
@@ 2. $(\mathbf{noq_{2}})$
    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{A}_{S}$.
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{B}$ is q-no-definable from $\mathbb{A}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
In the following, it shall be useful to observe that Lemma 2, Lemma 7Lemma 4, and Lemma 6 also hold in the non-oracular setting. Furthermore, the definition of quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ (Definition 3) does not change in the non-oracular setting.
The following result is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
@@
Let $\mathbb{A}$ be a finite structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$. Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{A}_{S})$.
@@
Henceforth in this proof, we shall denote $\mathbb{A}_{S}$ by $\mathbb{D}$ for the sake of readability.
@@
    Let $\mathbb{G}$ be a non-oracular q-definition for $\mathbb{A}$ and $S$. This means that every quantum homomorphism $\mathbb{R}\to_{q}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{qno}}\mathbb{D}$.
    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{qno}}\mathbb{A}$. We want to prove that $Q\in\operatorname{qPol^{no}}(\mathbb{D})$. Note that $Q$ is a quantum function $A^{n}\to_{q}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
    Fix then such a tuple $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})\in R^{\mathbb{D}^{n}}$, and choose a tuple $\mathbf{c}\in A^{r}\setminus S$. For every $i\in[n]$, consider the deterministic homomorphism $f_{i}:\mathbb{R}\to\mathbb{D}$ defined by $f_{i}(j)=b^{(j)}_{i}$ for each $j\in[r]$. In particular, $f_{i}$ is a quantum homomorphism. Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$ extending $f_{i}$. Consider now the tensor product
    $W=\bigotimes_{i\in[n]}W^{(i)}:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}^{n},$
    and compose it with $Q$ to obtain $Q\bullet W:\mathbb{G}\operatorname{\to_{qno}}\mathbb{A}$. Fix $\ell\in[r]$ and $a\in A$. Expanding the composition and the tensor product yields
@@
    Let $Q\colon\mathbb{R}\to_{q}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$. Since $Q$ satisfies $($$\mathbf{QH_{2}}$)), $Q$ must be classical, so by the spectral theorem we can decompose it as $Q=\bigoplus_{i\in[d]}h_{i}$, where $h_{i}\colon\mathbb{R}\to\mathbb{D}$ is a deterministic homomorphism for each $i\in[d]$. By the definition of $\mathbb{R}$, it must be that $(h_{i}(1),\dots,h_{i}(r))\in S$ for each $i\in[d]$. In other words, there exists a function $f:[d]\to[n]$ satisfying
    (11) $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$
    for each $i\in[d]$. Consider now the deterministic dictator polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$. Define
@@
    Note that $P$ is a classical (a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
@@
    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{qno}}\mathbb{A}$ be a non-oracular quantum homomorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$. Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuples $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures. Since $Q$ is a non-oracular quantum polymorphism of $\mathbb{A}$, by assumption it is also an $n$-ary non-oracular quantum polymorphism of $\mathbb{D}$. Hence, for each $\mathbf{c}\in A^{r}\setminus S$, we must have
@@
Corollary 42.
Two structures $\mathbb{A},\mathbb{B}$ satisfy $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}(\mathbb{B})$ if and only if $\mathbb{B}$ is q-no-definable from $\mathbb{A}$.
Unlike in the oracular setting, non-oracular q-gadgets are not strong enough to guarantee reductions. To that end, they need to be strengthened via a non-oracular commutativity gadget, which we now define.

### v2

@@ 9 Quantum polymorphisms of odd cycles
If $x=z$ the result is trivial, so we may assume that $x\neq z$. We know from Corollary 71 that there exist $d\in\mathbb{N}$ and $h_{1},\dots,h_{d}\in\operatorname{Hom}(\mathbb{C}_{m},\mathbb{C}_{m})$ such that $Q=\bigoplus_{i\in[d]}h_{i}$. Using eq. 1, we can write
$Q_{x,a}Q_{y,b}Q_{z,c}$ $=\left(\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}\right)\left(\sum_{\begin{subarray}{c}j\in[d]\\ h_{j}(y)=b\end{subarray}}\mathbf{e}_{j}\mathbf{e}_{j}^{*}\right)\left(\sum_{\begin{subarray}{c}k\in[d]\\ h_{k}(z)=c\end{subarray}}\mathbf{e}_{k}\mathbf{e}_{k}^{*}\right)=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(x)=a\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}.$
@@
$Q_{x,a}Q_{y,b}Q_{z,c}$ $=\sum_{\begin{subarray}{c}i\in[d]\\ h_{i}(y)=b\\ h_{i}(z)=c\end{subarray}}\mathbf{e}_{i}\mathbf{e}_{i}^{*}=Q_{y,b}Q_{z,c},$
@@
Observe that the non-contextuality of unary quantum polymorphisms of odd cycles—i.e., quantum homomorphisms $\mathbb{C}_{m}\to_{\operatorname{q}}\mathbb{C}_{m}$—is clear from Corollary 71. Hence, in order to prove Theorem 85, we need to extend non-contextuality to higher arities. Henceforth in this section, we fix some arity $n\geq 2$. Furthermore, whenever a plus or minus sign is applied to an element of $\mathbb{Z}_{m}$ or to a tuple of such elements, it shall be meant modulo $m$.
@@
Hence, $Q_{\mathbf{x}}=(Q_{\mathbf{y}}+Q_{\mathbf{y}-2\mathbf{v}})Q_{\mathbf{x}}$. Since projectors are self-adjoint, it follows that
@@
We now use the same argument after applying the automorphism of $\mathbb{C}_{m}^{n}$ given by $\mathbf{a}\mapsto\mathbf{a}+\mathbf{y}-\mathbf{x}$. We obtain
@@
In particular, we deduce that $Q_{\mathbf{y},a+2}Q_{\mathbf{x},a}$ is self-adjoint, so
@@
We can now finalise the proof of Theorem 85.
@@
$Q_{\mathbf{x},a}Q_{\mathbf{y},b}$ $=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$ (16)
@@
$=\sum_{\mathbf{c}\in\mathbb{Z}_{m}^{\delta-1}}Q_{\mathbf{x}_{\mathbf{a}}}Q_{\mathbf{y},b}Q_{\mathbf{x},a}Q_{\mathbf{z}_{1},c_{1}}Q_{\mathbf{z}_{2},c_{2}}\dots Q_{\mathbf{z}_{\delta-1}c_{\delta-1}}.$
@@
where the second equality is eq. 16. From this, we conclude that $Q_{\mathbf{x},a}$ and $Q_{\mathbf{y},b}$ commute, as required. ∎
The undecidability of $\operatorname{CSP_{q}}(\mathbb{C}_{m})$ is now a simple consequence of the machinery developed in Section 4 and Section 5.
Proof of Theorem 9.
$\mathbb{K}_{m}$ is pp-definable in $\mathbb{C}_{m}$ with the gadget consisting of a path of length m2m-2; hence, $\operatorname{Pol}(\mathbb{C}_{m})\to\operatorname{Pol}(\mathbb{K}_{m})\to\operatorname{Pol}(\operatorname{3SAT})$, and the result follows by combining Theorem 85 and Corollary 56. ∎
We conclude this section by proving that quantum homomorphisms between different odd cycles can be contextual.
Proposition 89.
There is a quantum homomorphism $\mathbb{C}_{7}\to_{\operatorname{q}}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space that is not in $\oplus{\operatorname{Hom}(\mathbb{C}_{7},\mathbb{C}_{5})}$.
@@
Consider for this proof $\mathbb{C}_{5}$ to have domain $\{0,1,2,5,6\}$ with edges between adjacent numbers, and $\mathbb{C}_{7}$ to have domain $[7]$. Consider the two classical endomorphisms $h,g\colon\mathbb{C}_{5}\to\mathbb{C}_{5}$ such that $h(2)=6,h(5)=0$ and $g(5)=0,g(2)=1$. Then $h\oplus g$ is a quantum homomorphism $Q\colon\mathbb{C}_{5}\to_{\operatorname{q}}\mathbb{C}_{5}$ over a 2-dimensional Hilbert space. Extend $Q$ by setting $Q_{3,a}=Q_{5,a}$ for $a\in\{0,1,2,5,6\}$. Choose a basis in which all the projectors are diagonal, and in this basis define $Q_{4,1}=\frac{1}{2}\begin{pmatrix}1&1\\ 1&1\end{pmatrix}$ and $Q_{4,6}=\frac{1}{2}\begin{pmatrix}1&-1\\ -1&1\end{pmatrix}$. Then $Q\colon\mathbb{C}_{7}\to_{\operatorname{q}}\mathbb{C}_{5}$ and $Q_{4,1},Q_{2,1}]\neq 0$ since $Q_{2,1}=\begin{pmatrix}0&0\\ 0&1\end{pmatrix}$. ∎
10 Non-oracular polymorphisms
In this section, we consider the non-oracular version of the results described in Section 4 and Section 5. A non-oracular quantum homomorphism between two $\sigma$-structures $\mathbb{A}$ and $\mathbb{B}$, denoted by $Q\colon\mathbb{A}\operatorname{\to_{\operatorname{qno}}}\mathbb{B}$, is a quantum function $Q\colon A\to_{\operatorname{q}}B$ satisfying the condition $($$\mathbf{QH_{1}}$)) but possibly not $($$\mathbf{QH_{2}}$)); see Section 3.2. The following is the non-oracular version of the quantum CSP.
Definition 90 (Non-oracular quantum CSP).
Fix a $\sigma$-structure $\mathbb{A}$. $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ is the following computational problem: Given as input a $\sigma$-structure $\mathbb{X}$, determine whether there exists a non-oracular quantum homomorphism $\mathbb{X}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$.
For CSPs parameterised by digraphs, non-oracular quantum homomorphisms capture perfect strategies for the variant of the 2-player CSP game where both cooperating players receive some variable ($x_{A}$ and $x_{B}$, respectively) from the verifier and reply with assignments to their variables with some values $a_{A}$ and $a_{B}$, respectively. The answers are deemed correct by the verifier if $a_{A}=a_{B}$ whenever $x_{A}=x_{B}$ (i.e., the game is synchronous), and $(a_{A},a_{B})$ is an edge in the template digraph whenever $(x_{A},x_{B})$ is an edge in the instance. The relation between this game and the type of quantum homomorphisms given above was first described in [59] (in the case of undirected graphs) as a generalisation of the quantum colouring game of [23], see also [58]. Various other works have considered this setting for specific types of binary constraint systems; see, e.g., the non-oracular strategies for Unique Projection Games [51] considered in [50].
We next define quantum polymorphisms in the non-oracular setting.
@@ 10 Non-oracular polymorphisms
Let $\mathbb{A}$ be a relational structure. A non-oracular quantum polymorphism is a non-oracular quantum homomorphism $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$. We denote by $\operatorname{qPol^{no}}(\mathbb{A})$ the set of non-oracular quantum polymorphisms of $\mathbb{A}$.
As the main result of this section, we shall prove that, just like in the oracular case Theorem 5, the Galois connection between classical polymorphism clones and pp-definitions [15, 38] can be lifted to a Galois connection for non-oracular quantum CSPs. As a consequence, we will prove the following result.
@@
Then $\operatorname{CSP_{q}}(\mathbb{B})$ reduces to $\operatorname{CSP_{q}^{no}}(\mathbb{A})$ in logspace.
It should be observed that, in this case, we obtain a reduction from the oracular version of the quantum CSP parameterised by $\mathbb{B}$ to the non-oracular version of the quantum CSP parameterised by $\mathbb{A}$. As we shall see, this is a consequence of how commutativity in the contexts is handled by the reduction.
@@
Let $\mathbb{A}$ be a $\sigma$-structure and let $S\subseteq A^{r}$ be a set of $r$-tuples of elements of $A$ for some $r\in\mathbb{N}$. Let $R$ be the $r$-ary symbol of the signature of $(A;S)$ (i.e., $R^{(A;S)}=S$), and consider the $\{R\}$-structure $\mathbb{R}$ having domain $[r]$ and relation $R^{\mathbb{R}}=\{(1,\dots,r)\}$. A non-oracular q-definition (in short, a q-no-definition) for $\mathbb{A}$ and $S$ is a $\sigma$-structure $\mathbb{G}$ with distinguished vertices $g_{1},\dots,g_{r}$ satisfying the following properties:
@@ 1. $(\mathbf{noq_{1}})$
    For every quantum homomorphism $Q\colon\mathbb{R}\to_{\operatorname{q}}(A;S)$, there exists a non-oracular quantum homomorphism $Q^{\prime}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ such that $Q^{\prime}_{g_{i},a}=Q_{i,a}$ for all $i\in[r]$ and all $a\in A$.
@@ 2. $(\mathbf{noq_{2}})$
    For every non-oracular quantum homomorphism $Q\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, the quantum function $Q^{\prime}$ defined by $Q^{\prime}_{i,a}:=Q_{g_{i},a}$ for all $i\in[r]$ is a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}(A;S)$.
@@
Let $\mathbb{A}$ and $\mathbb{B}$ be structures on the same domain and potentially different signatures. We say that $\mathbb{A}$ q-no-defines $\mathbb{B}$ if there exists a non-oracular q-definition for $\mathbb{A}$ and $R^{\mathbb{B}}$ for every relation $R^{\mathbb{B}}$ of $\mathbb{B}$.
In the following, it shall be useful to observe that Lemma 17, Lemma 24Lemma 19, and Lemma 21 also hold in the non-oracular setting. Furthermore, the definition of quantum closure of $\operatorname{Hom}(\mathbb{A},\mathbb{B})$ (Definition 18) does not change in the non-oracular setting. The next result is the non-oracular version of Theorem 40, and is the key to the quantisation of the relation/operation Galois connection in the non-oracular quantum setting.
@@
Let $\mathbb{A}$ be a relational structure and let $S\subseteq A^{r}$ for some $r\in\mathbb{N}$. Then there exists a non-oracular q-definition for $\mathbb{A}$ and $S$ if, and only if, $\operatorname{qPol^{no}}(\mathbb{A})\subseteq\operatorname{qPol^{no}}((A;S))$.
@@
Henceforth in this proof, we shall denote $(A;S)$ by $\mathbb{D}$ for the sake of readability.
@@
    Let $\mathbb{G}$ be a non-oracular q-definition for $\mathbb{A}$ and $S$. This means that every quantum homomorphism $\mathbb{R}\to_{\operatorname{q}}\mathbb{D}$ extends to a non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$, and every non-oracular quantum homomorphism $\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ restricts to a non-oracular quantum homomorphism $\mathbb{R}\operatorname{\to_{\operatorname{qno}}}\mathbb{D}$.
    Fix now a non-oracular quantum polymorphism $Q$ of $\mathbb{A}$; i.e., $Q\colon\mathbb{A}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$. We want to prove that $Q\in\operatorname{qPol^{no}}(\mathbb{D})$. Note that $Q$ is a quantum function $A^{n}\to_{\operatorname{q}}A$, so we only need to prove $($$\mathbf{QH_{1}}$)) applied to the relation of $\mathbb{D}^{n}$; i.e., to the relation $R^{\mathbb{D}^{n}}$ containing all $r$-tuples $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})$ where each $\mathbf{b}^{(i)}$ is an element of $A^{n}$, and for each $j\in[n]$ it holds that $(b^{(1)}_{j},\dots,b^{(r)}_{j})\in R^{\mathbb{D}}=S$.
    Fix then such a tuple $(\mathbf{b}^{(1)},\dots,\mathbf{b}^{(r)})\in R^{\mathbb{D}^{n}}$, and choose a tuple $\mathbf{c}\in A^{r}\setminus S$. For every $i\in[n]$, consider the deterministic homomorphism $f_{i}:\mathbb{R}\to\mathbb{D}$ defined by $f_{i}(j)=b^{(j)}_{i}$ for each $j\in[r]$. In particular, $f_{i}$ is a quantum homomorphism. Hence, ($(\mathbf{noq_{1}})$) implies that there exists a non-oracular quantum homomorphism $W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ extending $f_{i}$. Consider now the tensor product
    $W=\bigotimes_{i\in[n]}W^{(i)}\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}^{n},$
    and compose it with $Q$ to obtain $Q\bullet W\colon\mathbb{G}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$. Fix $\ell\in[r]$ and $a\in A$. Expanding the composition and the tensor product yields
@@
    Let $Q\colon\mathbb{R}\to_{\operatorname{q}}\mathbb{D}$ be a quantum homomorphism over some Hilbert space $H$, and let $d$ be the dimension of $H$. Since $Q$ satisfies $($$\mathbf{QH_{2}}$)), $Q$ must be classical, so by the spectral theorem we can decompose it as $Q=\bigoplus_{i\in[d]}h_{i}$, where $h_{i}\colon\mathbb{R}\to\mathbb{D}$ is a deterministic homomorphism for each $i\in[d]$. By the definition of $\mathbb{R}$, it must be that $(h_{i}(1),\dots,h_{i}(r))\in S$ for each $i\in[d]$. In other words, there exists a function $f:[d]\to[n]$ satisfying
    $(h_{i}(1),\dots,h_{i}(r))=\mathbf{a}^{(f(i))}$ (17)
    for each $i\in[d]$. Consider now the classical projection polymorphisms $\pi^{n}_{j}:\mathbb{A}^{n}\to\mathbb{A}$ defined by $\pi^{n}_{j}(\mathbf{a})=a_{j}$ for each $j\in[n]$ and each $\mathbf{a}\in A^{n}$. Define
@@
    Note that $P$ is a classical (so, a fortiori, non-oracular quantum) homomorphism from $\mathbb{A}^{n}$ to $\mathbb{A}$. Take now $\ell\in[r]$ and $a\in A$, and observe that
@@
    Finally, let $Q\colon\mathbb{\mathbb{A}}^{n}\operatorname{\to_{\operatorname{qno}}}\mathbb{A}$ be a non-oracular quantum polymorphism, and consider the quantum function $Q^{\prime}$ given by $Q^{\prime}_{\ell,a}=Q_{g_{\ell},a}$ for $\ell\in[r]$ and $a\in A$. Recall now that each $\mathbf{a}^{(i)}$ is an element of $S=R^{\mathbb{D}}$ and, thus, the tuple $(g_{1},\dots,g_{r})$ belongs to $R^{\mathbb{D}^{n}}$ by the definition of categorical power of structures. Since $Q$ is a non-oracular quantum polymorphism of $\mathbb{A}$, by assumption it is also an $n$-ary non-oracular quantum polymorphism of $\mathbb{D}$. Hence, for each $\mathbf{c}\in A^{r}\setminus S$, we must have
@@
See 13
Unlike in the oracular setting, non-oracular q-gadgets alone are not strong enough to guarantee reductions. To that end, they need to be strengthened via a non-oracular commutativity gadget, which we now define.